OpenStudy (canada907cat):

Identify and remove removable the discontinuities of 1) f(x)=sin(x+1)/tan (x+1)

OpenStudy (mathmate):

@Canada907cat Do you know where the discontinuity is, with the function as it is? Recall that 1. a discontinuity can be caused by the denominator becoming zero 2. tan(x)=sin(x)/cos(x).

OpenStudy (canada907cat):

I really don't know, this subject is completely new to me.

OpenStudy (amirrezas):

It has discontinuity in \[x=-1+\pi/2 +k\pi] All of them removable with putting the value of cos(x+1) instead of the f(x) in discontinuous points.

OpenStudy (canada907cat):

so would that mean that the discontinuity is at x=-1?

OpenStudy (vanesaretana):

..Wow

OpenStudy (amirrezas):

The discontinuity is not at x=-1, but It is at x=-1+\pi/2. discontinuity is not having one of this: 1)left limit 2)right limit 3)f(that point) In this case the f(x) does not exist because of denominator

OpenStudy (canada907cat):

Okay well I have to choose either x=-1 or x=0 for the discontinuity .

OpenStudy (amirrezas):

Its x=-1 because it makes the denominator zero.

OpenStudy (canada907cat):

Okay so the next part of the question is that can it be removed by making f(-1)=1 or can not be removed and based on what you just said I'm guessing it cannot be removed.

OpenStudy (amirrezas):

Yes it can be removed by setting f(-1)=1 because cos(x+1) is a continuous function which can be replaced in non discontinuous points of given f(x) cos(x+1)=cos(-1+1)=1 Therefore we have both left n right limit and function value at -1. The modified function is continuous in point x=-1

OpenStudy (canada907cat):

oh okay that makes sense.

OpenStudy (amirrezas):

Also you have to remember that condition for being continuous is that left and right limit are both equal to the f(x) in the point you are analyzing continuity.

OpenStudy (canada907cat):

okay great thank you so can you help me one more question that has the same concept as the first one?

OpenStudy (amirrezas):

Make a new question!

OpenStudy (canada907cat):

okay will do.